# Nakai-Moishezon criteria and the toric Thomas-Yau conjecture

### v1

@@ 5.2.2 Complete intersections
Thus, by (5.4), the phase inequality (3.2) in this case
@@ 5.2.3 Induction
$\int_{V_{2}\pitchfork\cdots\pitchfork V_{r}}e^{-\operatorname{i}[\omega_{0}]}\operatorname{ch}(L)$
$=(2\pi\operatorname{i})^{-n}k^{r-1}{\hat{k}}^{-n+1}\prod^{r}_{i=2}k_{i}\int_{\Gamma(\mathcal{S}(k\hat{k}L,\hat{k}L^{k}(k_{2}V_{2}),\ldots,k_{r}V_{r}))}e^{-W(k\hat{k}\omega_{0})/z}\Omega_{0}$
@@
$\int_{V_{1}\pitchfork\cdots\pitchfork V_{r}}e^{-\operatorname{i}[\omega_{0}]}\operatorname{ch}(L)$
$=-(2\pi\operatorname{i})^{-n}k^{r}{\hat{k}}^{-n+1}\prod^{r}_{i=1}k_{i}\int_{\Gamma(\mathcal{S}_{V_{1}\pitchfork\cdots\pitchfork V_{r}})}e^{-W(k\hat{k}\omega_{0})/z}\Omega_{0}$
@@
We can now choose our constants
@@ 6 Relation to Bridgeland stability on surfaces
In this Section we study Theorem 2.5 and its Corollaries 2.7, 2.9 from the viewpoint of Bridgeland stability, when $X$ is a weak del Pezzo surface. In particular we prove Proposition 1.8.
Fix a Kähler class $[\omega_{0}]$. There is well-known Bridgeland stability condition $\sigma_{[\omega_{0}]}$ on $D^{b}(X)$, with central charge given by
@@
Suppose that we are in the situation of Theorem 2.5, on a weak del Pezzo $X$. In the proof of Theorem 2.5, Section 5.2.1, we constructed objects $\mathcal{S}(kL,k_{V}V)\in D^{b}(X)$, with morphisms
@@
Assume now that $L^{\vee}$ is semiample. Then, we have
@@
since $L^{k}$ is a seminegative line bundle and so $\mu_{[\omega_{0}]}\leq 0$ for all $[\omega_{0}]$. On the other hand,
@@
This argument shows that, when $L^{\vee}$ is semiample, the morphism (6.1) is actually a morphism in the abelian category $\operatorname{Coh}^{\sharp}(X)$.
@@
Let $X$ be a toric surface. Fix a Kähler class $[\omega_{0}]$. Suppose $L^{\vee}$ is semiample and fix $k>0$ sufficiently large. Then, the dHYM equation is solvable on $L^{\vee}$ if, and only if, the object $L^{k}[1]$ is Bridgeland stable with respect to the stability condition
@@
When $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ the converse also holds.
@@
Suppose that $L^{k}[1]$ is stable with respect to $\sigma_{k[\omega_{0}]}$. Since we are assuming that $L^{\vee}$ is semiample, the morphisms (6.1) induce exact triangles of objects contained in the heart $\operatorname{Coh}^{\sharp}(X)$. Thus, by the definition of Bridgeland stability, the phase inequalities for central charges (6.2) must hold for all toric divisors $V$. Therefore, by the proof of Theorem 2.5, Section 5.2.1, the line bundle $L^{k}$ is dHYM positive with respect to $k[\omega_{0}]$. So the same holds for $L$ with respect to $[\omega_{0}]$, and the dHYM equation with respect to $[\omega_{0}]$ is solvable on $L$ and $L^{\vee}$.
In the other direction, Collins and Shi [11], Section 4 (building on the results of Arcara-Miles [3]), show that on $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$, for any $[\omega_{0}]$, if the dHYM equation is solvable on a line bundle, non necessarily semiample, then the line bundle is Bridgeland stable with respect to $\sigma_{[\omega_{0}]}$. On the other hand, the dHYM equation is solvable on $L^{\vee}$ with respect to $[\omega_{0}]$ iff it is solvable on $L^{k}$ with respect to $k[\omega_{0}]$, so, if dHYM is solvable on $L^{\vee}$ with respect to $[\omega_{0}]$, then $L^{k}$ is stable with respect to the stability condition $\sigma_{k[\omega_{0}]}$, and the same holds for the shift $L^{k}[1]$ (by the definition of Bridgeland stability). ∎
@@
Note that $L$ is semiample, so does not fit our conventions, although probably the example can be modified.
@@
Firstly, note that the special Lagrangian condition for the shifted section $\tilde{\mathcal{L}}$ coincides with the dHYM equation for the the mirror line bundle $L^{k}$ with respect to $k[\omega_{0}]$. This is solvable iff it is solvable for $L^{\vee}$ with respect to $[\omega_{0}]$. Since $L^{\vee}$ is ample by assumption, by Proposition 6.1, this holds if the object $L^{k}[1]$ is Bridgeland stable with respect to the stability condition $\sigma_{k[\omega_{0}]}$, i.e. if $\tilde{\mathcal{L}}$ is Bridgeland stable in the sense of Definition 1.5. The converse also holds on $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ by Proposition 6.1.

### v2

@@ 5.2.2 Complete intersections
Suppose that $[\omega_{0}]$ is generic with respect to $V_{1}$, $V_{2}$ in the sense that it does not lie in the uion of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$, $V_{1}$, $V_{2}$.
Then, by (5.4), the phase inequality (3.2) in this case
@@ 5.2.3 Induction
$\int_{V_{2}\pitchfork\cdots\pitchfork V_{r}}e^{-\operatorname{i}[\omega_{0}]}\operatorname{ch}(L)$ $=(2\pi\operatorname{i})^{-n}k^{r-1}{\hat{k}}^{-n+1}\prod^{r}_{i=2}k_{i}\int_{\Gamma(\mathcal{S}(k\hat{k}L,\hat{k}L^{k}(k_{2}V_{2}),\ldots,k_{r}V_{r}))}e^{-W(k\hat{k}\omega_{0})/z}\Omega_{0}$
@@
$\int_{V_{1}\pitchfork\cdots\pitchfork V_{r}}e^{-\operatorname{i}[\omega_{0}]}\operatorname{ch}(L)$ $=-(2\pi\operatorname{i})^{-n}k^{r}{\hat{k}}^{-n+1}\prod^{r}_{i=1}k_{i}\int_{\Gamma(\mathcal{S}_{V_{1}\pitchfork\cdots\pitchfork V_{r}})}e^{-W(k\hat{k}\omega_{0})/z}\Omega_{0}$
@@
Suppose that $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$ and the classes of all toric submanifolds.
We can then choose our constants
@@ 6 Relation to Bridgeland stability on surfaces
In this Section we study Theorem 2.5 and its Corollaries 2.7, 2.9 from the viewpoint of Bridgeland stability, when $X$ is a weak del Pezzo surface. In particular we prove Theorem 1.9.
Fix a Kähler class $[\omega_{0}]$. There is well-known Bridgeland stability condition $\sigma_{[\omega_{0}]}$ on $D^{b}(X)$,
$\sigma_{[\omega_{0}]}=\left(\operatorname{Coh}^{\sharp}(X),Z_{[\omega_{0}]}\right),$
with central charge given by
@@
Suppose that we are in the situation of Theorem 2.5, on a weak del Pezzo $X$. In the proof of Theorem 2.5, Section 5.2.1 and Remark 5.5, we constructed objects $\mathcal{S}(kL,k_{V}V)\in D^{b}(X)$, with morphisms
@@
Assume now that $L^{\vee}$ is ample. Then, we have
@@
since $L^{k}$ is a negative line bundle and so $\mu_{[\omega_{0}]}\leq 0$ for all $[\omega_{0}]$. On the other hand,
@@
This argument shows that, when $L^{\vee}$ is positive, the morphism (6.1) is actually a morphism in the abelian category $\operatorname{Coh}^{\sharp}(X)$.
For our applications we also need to recall a conjecture of Arcara and Miles.
Conjecture 6.1 ([3], Conjecture 1).
Let $X$ be a smooth projective surface with a fixed Kähler class $[\omega_{0}]$ and a line bundle $L$. Then,
  1. $(i)$
    if $L$ is unstable with respect to $\sigma_{[\omega_{0}]}$, then it is destabilised by $L(-C)$, where $C$ is a curve of negatibe self-intersection;
  2. $(ii)$
    if $L[1]$ is unstable with respect to $\sigma_{[\omega_{0}]}$, then it is destabilised by $L(C)|_{C}$, where $C$ is a curve of negatibe self-intersection.
Note that [3], Conjecture 1 actually allows more general (divisorial) stability conditions. The conjecture is known for $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ (see [3]) and for $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$ (see the recent work [31]).
@@ 2. $(ii)$
Let $X$ be a toric surface. Fix a Kähler class $[\omega_{0}]$ and a positive line bundle $L^{\vee}$. Fix $k>0$ sufficiently large. Suppose $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$. Then, the dHYM equation is solvable on $L^{\vee}$ with respect to $[\omega_{0}]$ if, and only if, the object $L^{k}[1]$ is Bridgeland stable with respect to the stability condition
@@
The converse also holds, conditionally on the Arcara-Miles conjecture 6.1, which is known when $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ or $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$.
@@
Suppose that $L^{k}[1]$ is stable with respect to $\sigma_{k[\omega_{0}]}$. Since we are assuming that $L^{\vee}$ is positive, the morphisms (6.1) induce exact triangles of objects contained in the heart $\operatorname{Coh}^{\sharp}(X)$. Thus, by the definition of Bridgeland stability, the phase inequalities for central charges (6.2) must hold for all toric divisors $V$. Therefore, under our assumptions, by the proof of Theorem 2.5, Section 5.2.1 and Remark 5.5, the line bundle $L^{k}$ is dHYM positive with respect to $k[\omega_{0}]$. So the same holds for $L$ with respect to $[\omega_{0}]$, and the dHYM equation with respect to $[\omega_{0}]$ is solvable on $L$ and $L^{\vee}$.
Conversely, suppose that the dHYM equation is solvable on the positive line bundle $L^{\vee}$ with respect to $[\omega_{0}]$. Thus, it is also solvable on the negative line bundle $L$. Following Collins-Shi [11], proof of Theorem 4.15, we write the corresponding dHYM positivity condition as
$C\cdot\operatorname{ch}_{1}(L)>\frac{\operatorname{ch}_{2}(L)-\frac{1}{2}[\omega_{0}]^{2}}{\operatorname{ch}_{1}(L)\cdot[\omega_{0}]}(C\cdot[\omega_{0}]),$
for all curves $C\subset X$. As we know, this must be invariant under our rescaling, i.e. equivalent to
$C\cdot\operatorname{ch}_{1}(L^{k})>\frac{\operatorname{ch}_{2}(L^{k})-\frac{1}{2}[k\omega_{0}]^{2}}{\operatorname{ch}_{1}(L^{k})\cdot[k\omega_{0}]}(C\cdot[k\omega_{0}]).$
On the other hand, assuming the Arcara-Miles conjecture 6.1, $L^{k}$ is only potentially destabilised with respect to $\sigma_{k[\omega_{0}]}$ by the objects $L^{k}(-C)$ where $C$ is a curve with $C^{2}<0$. The condition that such objects do not destabilise is given by
$\frac{\operatorname{ch}_{2}(L^{k})-C\cdot\operatorname{ch}_{1}(L^{k})+\frac{1}{2}C^{2}-\frac{1}{2}[k\omega_{0}]^{2}}{\operatorname{ch}_{1}(L^{k})\cdot[k\omega_{0}]-C\cdot[k\omega_{0}]}<\frac{\operatorname{ch}_{2}(L^{k})-\frac{1}{2}[k\omega_{0}]^{2}}{\operatorname{ch}_{1}(L^{k})\cdot[k\omega_{0}]}.$
Following again [11], proof of Theorem 4.15, we note that this is equivalent to the condition
$C\cdot\operatorname{ch}_{1}(L^{k})-\frac{1}{2}C^{2}>\frac{\operatorname{ch}_{2}(L^{k})-\frac{1}{2}[k\omega_{0}]^{2}}{\operatorname{ch}_{1}(L^{k})\cdot[k\omega_{0}]}(C\cdot[k\omega_{0}]).$
Since $C^{2}<0$, this shows that if the dHYM equation is solvable on $L^{\vee}$, then the line bundle $L^{k}$ is stable with respect to $\sigma_{k[\omega_{0}]}$.
Finally, since $L$ is negative, the object $L^{k}[1]$ lies in the heart $\operatorname{Coh}^{\sharp}(X)$ of $\sigma_{k[\omega_{0}]}$. Thus it is stable with respect to $\sigma_{k[\omega_{0}]}$ iff its shift $L^{k}$ is, by the definition of Bridgeland stability condition. It follows that if the dHYM equation is solvable on $L^{\vee}$ then $L^{k}[1]$ is stable with respect to $\sigma_{k[\omega_{0}]}$. ∎
@@
Note that $L$ is positive, so does not fit our conventions, although probably the example can be modified.
@@
Firstly, note that the special Lagrangian condition for the shifted section $\tilde{\mathcal{L}}$ coincides with the dHYM equation for the the mirror line bundle $L^{k}$ with respect to $k[\omega_{0}]$. This is solvable iff it is solvable for $L^{\vee}$ with respect to $[\omega_{0}]$. Since $L^{\vee}$ is ample by assumption, by Proposition 6.2, this holds if the object $L^{k}[1]$ is Bridgeland stable with respect to the stability condition $\sigma_{k[\omega_{0}]}$, i.e. if $\tilde{\mathcal{L}}$ is Bridgeland stable in the sense of Definition 1.6. The converse also holds conditionally on the Arcara-Miles conjecture 6.1, by Proposition 6.2, and so for $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ or $X=\operatorname{Bl}_{p,q}\mathbb{P}^{2}$.
6.1 The unstable case
In this Section we prove Theorem 1.12. The proof is an application of a result on weak solutions of the dHYM equation of surfaces due to Datar-Mete-Song [15], which we briefly recall.
Let $(X,[\omega_{0}],[\alpha_{0}])$ denote a surface with Kähler class and a $(1,1)$-class $[\alpha_{0}]$. The reference [15], Section 1.3 introduces a minimal angle $\theta_{min}\in(0,\pi)$ determined by
$\cot\theta_{min}:=\sup_{D}\left\{\frac{([\alpha_{0}]-D)^{2}-[\omega_{0}]^{2}}{2([\alpha_{0}]-D)\cdot[\omega_{0}]},\,([\alpha_{0}]-D)\cdot[\omega_{0}]>0\right\},$
where the supremum is taken over effective $\mathbb{R}$-divisors. It is shown that the supremum is in fact achieved by some effective $\mathbb{R}$-divisor $D$, so we have
$\cot\theta_{min}=\frac{([\alpha_{0}]-D)^{2}-[\omega_{0}]^{2}}{2([\alpha_{0}]-D)\cdot[\omega_{0}]}.$
The mimimal angle $\theta_{min}$ should be compared to the topological angle $\varphi=\varphi(X,c_{1}(L),[\omega_{0}])$ appearing in the Nakai-Moishezon criterion (2.3). One shows the inequality
$\cot\theta_{min}\geq\cot\varphi(X,c_{1}(L),[\omega_{0}]),$
such that equality holds if and only if $(X,[\omega_{0}],c_{1}(L))$ satisfies the dHYM semipositivity condition
$\operatorname{Re}\left(c_{1}(L)+\operatorname{i}[\omega_{0}]\right)^{k}\cdot\gamma^{m-k}\cdot Z\geq\cot(\varphi)\operatorname{Im}\left(c_{1}(L)+\operatorname{i}[\omega_{0}]\right)^{k}\cdot\gamma^{m-k}\cdot Z$ (6.3)
where $Z$ ranges though $m$-dimensional analytic subvarieties, $\gamma$ ranges through Kähler classes, and $k=1,\ldots,m$.
Theorem 6.4 ([15], Theorem 1.10).
Suppose $[\alpha_{0}]$, $[\omega_{0}]$ satisfy $[\alpha_{0}]\cdot[\omega_{0}]>0$. Then, for any Kähler form $\omega\in[\omega_{0}]$, there exists a unique closed current $\mathcal{T}\in[\alpha_{0}]$ such that
$\operatorname{Re}\langle(\mathcal{T}+\operatorname{i}\omega)^{2}\rangle=\cot(\theta_{min})\operatorname{Im}\langle(\mathcal{T}+\operatorname{i}\omega)^{2}\rangle,\,\mathcal{T}\geq\cot(\theta_{min})\omega.$
The equation holds in the sense of measures, where delimited-⟨$\langle$-$\rangle$ - ⟩ denotes the non-pluripolar part.
Moreover we have
$\cot\theta_{min}\geq\cot\varphi(X,c_{1}(L),[\omega_{0}]),$
with equality iff $(X,[\omega_{0}],L)$ satisfies the dHYM semipositivity condition (6.3).
6.1.1 Proof of Theorem 1.12
Suppose the SYZ Lagrangian $\tilde{\mathcal{L}}$ is Bridgeland unstable. By definition, this means that the mirror shifted line bundle $L^{k}[1]$ is unstable (i.e. not semistable) with respect to $\sigma_{k[\omega_{0}]}$. Under our assumptions, $L^{k}[1]$ is an element of $\operatorname{Coh}^{\sharp}(X)$, so the instability of $L^{k}[1]$ implies the instability of $L^{k}$. As in the proof of Proposition 6.2, assuming the Arcara-Miles conjecture 6.1, this means that there exists a curve $C\subset X$, with $C^{2}<0$, such that
$C\cdot\operatorname{ch}_{1}(L^{k})-\frac{1}{2}C^{2}<\frac{\operatorname{ch}_{2}(L^{k})-\frac{1}{2}[k\omega_{0}]^{2}}{\operatorname{ch}_{1}(L^{k})\cdot[k\omega_{0}]}(C\cdot[k\omega_{0}]).$
Then we must have
$C\cdot\operatorname{ch}_{1}(L)<\frac{\operatorname{ch}_{2}(L)-\frac{1}{2}[\omega_{0}]^{2}}{\operatorname{ch}_{1}(L)\cdot[\omega_{0}]}(C\cdot[\omega_{0}]),$
which implies that the dHYM semipositivity condition (6.3) cannot hold.
Thus, according to Theorem 6.4, there exists a nontrivial effective $\mathbb{R}$-divisor $D$, such that
$\cot\theta_{min}=\frac{(-c_{1}(L)-D)^{2}-[\omega_{0}]^{2}}{2(-c_{1}(L)-D)\cdot[\omega_{0}]},\,(-c_{1}(L)-D)\cdot[\omega_{0}]>0$
with
$\cot\theta_{min}>\cot\varphi(X,-c_{1}(L),[\omega_{0}]).$
Note that since $D$ attains the supremum in the definition of $\cot\theta_{min}$, we must have
$\cot\theta_{min}=\cot\varphi(X,-L-D,[\omega_{0}]).$
Then, it follows from Theorem 6.4 again that there exists a unique weak solution of the dHYM equation (2.1), with the correct topological angle $e^{\operatorname{i}\phi}$, i.e. a unique closed current $\mathcal{T}_{1}\in c_{1}(-L-D)$ which satisfies the equation
$\operatorname{Re}\langle(\mathcal{T}_{1}+\operatorname{i}\omega_{0})^{2}\rangle=\cot(\varphi(X,-L-D,[\omega_{0}])\operatorname{Im}\langle(\mathcal{T}_{1}+\operatorname{i}\omega_{0})^{2}\rangle$ (6.4)
in the sense of measures, where delimited-⟨$\langle$-$\rangle$ - ⟩ denotes the non-pluripolar part, and such that
$\mathcal{T}_{1}-\cot(\varphi(X,-L-D,[\omega_{0}])\omega_{0}\geq 0.$
Suppose now that $D$ is in fact a $\mathbb{Q}$-divisor. Then, possibly by taking $k$ even larger, we can assume that $-kL-kD$ is a genuine line bundle. The morphism in $D^{b}(X)$
$L^{k}[1]\to L^{k}(kD)[1]$
induces a morphism in $D\!\operatorname{FS}(\mathcal{Y}_{q_{k}},W(k\omega_{0}))$,
~~$\tilde{\mathcal{L}}\to\tilde{\mathcal{L}}_{D},$
where we set ~$\tilde{\mathcal{L}}_{D}:=\mathcal{L}_{D}[1]$ for $\mathcal{L}_{D}$ a SYZ transform of $L^{k}(kD)$ with respect to $k\omega_{0}$.
Since $\mathcal{T}_{1}$ is the curvature of a singular Hermitian metric on $-L-D$, the equality of measures (6.4), which is invariant under rescaling $\mathcal{T}_{1}$ and $\omega_{0}$ by $k>0$, says that ~$\tilde{\mathcal{L}}_{D}$ is Hamiltonian isotopic, in a weak sense, to a (shifted) Lagrangian current section which is a weak solution of the special Lagrangian equation with phase angle $\theta_{min}$.
Remark 6.5.
Note that, by Theorem 6.4, the class $c_{1}(-L-D)$ satisfies the dHYM semipositivity condition (6.3) with respect to $[\omega_{0}]$. This implies that the Nakai-Moishezon criterion (2.3) holds weakly, i.e. for all proper irreducible subvarieties $V\subset X$ we have
$\int_{V}\operatorname{Re}(\operatorname{i}\omega_{0}+c_{1}(-L-D))^{\dim V}-\cot(\varphi)\operatorname{Im}(\operatorname{i}\omega_{0}+c_{1}(-L-D))^{\dim V}\geq 0.$ (6.5)
As in the proof of Proposition 6.2, the numerical semipositivity condition (6.5) implies that $L^{k}(kD)[1]$ is in fact stable with respect to $\sigma_{k[\omega_{0}]}$. However, in general, this does not automatically show that ~$\tilde{\mathcal{L}}_{D}$ is Hamiltonian isotopic to a smooth special Lagrangian (as in Theorem 1.9). For this, $[\omega_{0}]$ would need to be generic with respect to $c_{1}(L^{k}(kD))$, but the divisor $D$ in turn depends on $[\omega_{0}]$.
Example 6.6.
For the sake of this example we change our sign convention to fit with [15], Theorem 1.13. Consider $X=\operatorname{Bl}_{p}\mathbb{P}^{2}$ endowed with the Kähler class $[\omega_{0}]=b[H]-[E]$ and the line bundle $L^{\prime}=pH-qE$ (with $c_{1}(L^{\prime})\cdot[\omega_{0}]>0$). If $q<\cot(\varphi(X,L^{\prime},[\omega_{0}]))$, then $L^{\prime}$ does not support a smooth dHYM solution, we have $\cot(\theta_{min})>\cot(\varphi(X,L^{\prime},[\omega_{0}]))$, and there is a unique closed current $\mathcal{T}\in c_{1}(L^{\prime})$ such that
$\operatorname{Re}\langle(\mathcal{T}+\operatorname{i}\omega_{0})^{2}\rangle=\cot(\theta_{min})\operatorname{Im}\langle(\mathcal{T}+\operatorname{i}\omega_{0})^{2}\rangle,\,\mathcal{T}\geq\cot(\theta_{min})[\omega_{0}].$
According to [15] Theorem 1.13 (3), we have in this case
$\mathcal{T}=\chi+(\cot(\theta_{min})-q)[E],$
where $\chi$ is a $(1,1)$-form with bounded local potentials, solving the equation
$\operatorname{Re}(\chi+\operatorname{i}\omega_{0})^{2}=\cot(\theta_{min})\operatorname{Im}(\chi+\operatorname{i}\omega_{0})^{2}$
in the sense of Bedford-Taylor on $X\setminus E$.
Moreover, we have
$\cot(\theta_{min})=\cot\varphi(X,L^{\prime}(-D),[\omega_{0}])$
for some effective $\mathbb{R}$-divisor given by $D=tE$, for some $t>0$, realising the supremum
$\sup_{t}\left\{\frac{(c_{1}(L^{\prime})-tE)^{2}-[\omega_{0}]^{2}}{2(c_{1}(L^{\prime})-tE)\cdot[\omega_{0}]},\,(c_{1}(L^{\prime})-tE)\cdot[\omega_{0}]>0\right\}.$
It follows from [15] Theorem 1.13 (2) that the unique closed current $\mathcal{T}_{1}\in c_{1}(L^{\prime}(-D))$ such that
$\operatorname{Re}\langle(\mathcal{T}_{1}+\operatorname{i}\omega_{0})^{2}\rangle=\cot\varphi(X,L^{\prime}(-D),[\omega_{0}])\operatorname{Im}\langle(\mathcal{T}_{1}+\operatorname{i}\omega_{0})^{2}\rangle,\,\mathcal{T}_{1}\geq\cot\varphi(X,L^{\prime}(-D),[\omega_{0}])[\omega_{0}],$
is actually more regular. We have in fact $\mathcal{T}_{1}=\chi_{1}$, where $\chi_{1}$ is a $(1,1)$-form with bounded local potentials, satisfying the equation
$\operatorname{Re}(\chi_{1}+\operatorname{i}\omega_{0})^{2}=\cot\varphi(X,L^{\prime}(-D),[\omega_{0}])\operatorname{Im}(\chi_{1}+\operatorname{i}\omega_{0})^{2}$
in the sense of Bedford-Taylor on all $X$.